# GLINT-RU: Gated Lightweight Intelligent Recurrent Units for Sequential Recommender Systems

### v1

@@ 3.6. Complexity Analysis
In this subsection, we will analyze the time complexity of our GLINT-RU, and illustrate why GLINT-RU has inherent superiority over other popular SRS models in model efficiency.
Assuming that the sequence length is set as , the embedding size is and the kernel sizes for GLINT-RU is , the time complexity of GLINT-RU is . The complexity is calculated throughout the network, from the embedding layer to the prediction layer. GLINT-RU is more efficient than other models in the following aspects: Firstly, our model shows significantly low time complexity, as GLINT-RU is a highly paralleled mixed network with only one layer to achieve high performance. Our framework utilizes paralleled expert networks, and employs efficient GRU module to capture long-term dependencies, outperforming all the other models in theoretical complexity. Secondly, traditional transformer-based model SASRec (Kang and McAuley, 2018) suffer from large computational complexity, especially When the sequence length is large. LinRec (Liu et al., 2023a) changes the dot product of attention mechanism, substantially reduce the computational cost. However, Linrec still requires stacked transformers to achieve high performance. These transformer-based models require at least two layers to achieve satisfactory accuracy, while our GLINT-RU achieves outstanding performance with only one layer. Thirdly, Mamba4Rec utilizes a hardware-aware mechanism to accelerate the training and inference speed (Gu and Dao, 2023), but its improvement in recommendation performance and efficiency is not stable, which will be further verified via our experiments.
@@ 4. EXPERIMENTS
In this section, we conduct extensive experiments to show the effectiveness and efficiency of our GLINT-RU Framework. After we introduce our implementation detail, the experiment results will be analyzed in details. The experiments in this section are set to answer the following research questions:
@@
RQ3: How do dense selective GRU and linear attention mechanism contribute to GLINT-RU?
@@ 4.1. Datasets and Evaluation Metrics
We evaluate GLINT-RU based on three benchmark datasets ML-1M 111https://grouplens.org/datasets/movielens/, Amazon-Beauty and Amazon video Games 222https://cseweb.ucsd.edu/ jmcauley/datasets.html#amazon_reviews.
Statistical information of the three datasets is shown in Table 1. We adopt Recall, Mean Reciprocal Rank (MRR), and Normalized Discounted Cumulative Gain (NDCG) as the evaluation metrics for our experiments. The interactions are grouped by users chronologically, and the datasets are split by the leave-one-out strategy (Liu et al., 2023a).
@@ 4.2. Baselines
To demonstrate the effectiveness and efficiency of GLINT-RU, we make comparisons between GLINT-RU and state-of-the-art effective and efficient RNN-, attention-, MLP- and SSM-based baselines.
Traditional Models (1) GRU4Rec (Hidasi et al., 2015): utilizes GRUs to capture sequential dependencies within user interaction data for session-based recommendations. (2) BERT4Rec (Sun et al., 2019): adapts the Bidirectional Encoder Representations from Transformers (BERT) architecture to model user behaviors for personalized recommendation. (3) SASRec (Kang and McAuley, 2018): captures long-term and short-term user preferences by applying a multi-head attention mechanism to learn representations of user interaction sequences.
Efficient Models (1) LinRec (Liu et al., 2023a): reduces the computational costs substantially by changing the dot product of attention mechanism in the transformer-based models. We select SASRec as the backbone of LinRec. (2) SMLP4Rec (Gao et al., 2024) uses a tri-directional fusion scheme to learn correlations on sequence, channel and feature dimensions efficiently. (3) Mamba4Rec (Liu et al., 2024): explores the potential of selective SSMs for efficient sequential recommendation.
@@
Models | ML-1M | Amazon-Beauty | Amazon-Video-Games | ||||||
@@ 4.3. Implementation
In this subsection, we introduce the implementation details of the GLINT-RU. We use Adam optimizer (Kingma and Ba, 2014) with the learning rate 0.001 for our training process. Both the train and evaluation batch size are set as 2048. The hidden size is set as 128 for ML-1M, and 64 for Amazon Beauty and Video Games. As is shown in Table 1, the average length of ML-1M , Beauty and Video Games are 165, 8.88 and 9.54, so we set the maximum sequence length as 200 for ML-1M, and 100 for the two amazon datasets. We adopt the dropout rate 0.5 for amazon datasets considering their high level of sparsity, compared with 0.2 for ML-1M. Other implementation details follow the settings of original papers (Liu et al., 2024; Gao et al., 2024; Liu et al., 2023a).
@@ 4.4. Overall Performance (RQ1)
In this subsection, we compare the performance of GLINT-RU with both traditional recommendation frameworks and state-of-the-art efficient models. The results, as shown in Table 2, demonstrate the effectiveness of GLINT-RU on metrics Recall@10, MRR@10, and NDCG@10 in bold. According to the above table, it is evident that GLINT-RU defeats all the selected transformer-, RNN-, MLP- and SSM-based baselines. We improve the performance upper bound of efficient recommendation models by .
Traditional RNN-based models, like GRU4Rec, might have difficulty in dealing with complex user behaviors. Although it can capture the long-term dependencies from long sequences, it struggles to learn effectively from extremely sparse datasets. Traditional attention-based methods like SASRec have great performance on the three datasets, but it is quite inefficient due to the high computational complexity of the attention mechanism.
Efficient model LinRec changes the softmax operation, and takes attention scores from more positions into consideration, improving the performance on long-term sequential recommendation tasks compared with its backbone SASRec. Both SMLP4Rec and Mamba4Rec achieve impressive performance on the three datasets. However, Mamba4Rec demonstrates enhanced proficiency in modeling long-term sequence, but exhibits low performance in relatively short sequences. Conversely, the SMLP4Rec shows superior performance in short sequence recommendations while being less effective with longer sequences. In addition, SMLP4Rec requires additional features to enhance its performance. Uniquely, GLINT-RU integrates the advantages of linear attention mechanism and dense selective GRU module, adaptively extracting dependencies and most relevant behaviors from users. The gate mechanism employed in GLINT-RU substantially enhances its ability to filter the information based on the dynamic data environment and mix the experts according to the data adaptively.
@@
ML-1M | GLINT-RU | 31ms | 86s/epoch | 8.81G |
@@
SASRec | 55ms | 172s/epoch | 21.51GB | |
BERT4Rec | 88ms | 285s/epoch | 21.51GB | |
LinRec | 37ms | 101s/epoch | 11.67GB | |
SMLP4REc | 51ms | 151s/epoch | 16.13GB | |
@@
Beauty | GLINT-RU | 278ms | 3.8s/epoch | 2.62G |
@@
SASRec | 444ms | 8.1s/epoch | 7.67GB | |
BERT4Rec | 1372ms | 13s/epoch | 11.69GB | |
LinRec | 340ms | 5.6s/epoch | 4.14GB | |
SMLP4Rec | 361ms | 5.3s/epoch | 2.95GB | |
@@
Video Games | GLINT-RU | 247ms | 4.5s/epoch | 2.49G |
@@
SASRec | 406ms | 9.6s/epoch | 7.65GB | |
BERT4Rec | 1290ms | 15s/epoch | 10.98GB | |
LinRec | 327ms | 6.6s/epoch | 4.13GB | |
SMLP4Rec | 389ms | 9.1s/epoch | 3.38GB | |
@@
Inference time (ms) of each mini-batch, training time (s/epoch) and GPU memory (GB) of GLINT-RU and other baseline models. The best results are bolded, and the second best results are underlined.
@@ 4.5. Efficiency Comparison (RQ2)
In this subsection, we analyze the efficiency of GLINT-RU and state-of-the-art sequential recommendation models. We evaluate the model efficiency according to their inference time of each mini-batch, training time and GPU memory occupation.
The results, shown in Table 3, provide several valuable insights: Firstly, by utilizing the efficient Dense GRU module and linear attention module, GLINT-RU dramatically reduces the training time and inference time, improving the training inference time by compared with most efficient recommendation baseline models. In addition, due to the low computational cost of GRU and linear attention mechanism, GLINT-RU exhibits minimal GPU memory consumption, which is comparable to the state-of-the-art SSM-based efficient model Mamba4Rec. In Section 3.6, we demonstrate that the GLINT-RU exhibits low theoretical computational complexity, as we employs the parallel networks and efficient GRU as core component for our recommender system. The theoretical analysis has been verified by the results in Table 3.
@@
4.6. Ablation Study (RQ3)
In this subsection, we will analyze the efficacy of essential components in GLINT-RU architecture. We conduct the ablation study on the ML-1M dataset, and the results are outlined in Table 4, providing insightful observations.
The results verify the essential role of the dense selective GRU module, as the performance of the model will dramatically decrease without the GRU module. This reveals the insights that the gated GRU module effectively capture the long-term dependencies of user-item interactions. Linear attention mechanism supplements the information of relevant items in the sequence. As is shown in Table 4, it improves the performance of GLINT-RU to some extent. Adding a temporal convolution layer incorporates context information from adjacent items, resulting in an enhancement in model performance. In addition, the gated MLP block plays a similar role as the feed-forward network in our framework, which filters complex information from the expert mixing block. It is noteworthy that even without the gated MLP block our framework still outperforms all the state-of-the art efficient models, demonstrating its inherent remarkable superiority for sequential recommendation tasks. After we remove the gated MLP, the GPU memory occupation of GLINT-RU becomes 7.63GB, less than Mamba4Rec shown in Table 3, and the inference time will be reduced to 241ms. We name this framework as ”Light GLINT-RU”, which is more applicable to resource constrained scenarios.
@@ 4.7. Parameter Analysis
In this subsection, we analyze the impact of the crucial hyperparameter kernel size in GLINT-RU, which determines the amount of information of each hidden state. We discuss the performance and efficiency of GLINT-RU with varied , providing valuable insights of the stability and robustness of GLINT-RU.
We conduct the parameter analysis on the dataset Amazon Beauty, and the result is shown in Figure 3.
The first three subfigures display the model performance of GLINT-RU with different kernel sizes. Kernel sizes that are larger than 1 enable the aggregation of the information from adjacent items in each hidden state, enhancing the model performance. Additionally, our model exhibits stable and high performance, providing a wide range of choices of kernel sizes. This finding indicates the robustness of GLINT-RU framework. The accuracy of the recommendation system is observed to be positively correlated with the kernel size . This enhancement in performance can be attributed to the fact that a larger kernel size aggregates information from more items, thereby learning a more extensive context into the hidden state. However, as the kernel size continues to expand, dense selective GRU might incorporates irrelevant data into the output state, which might lead to a marginal decline in the accuracy. Mamba4Rec as a novel efficient SSM-based model, also employs a temporal convolution layer in the model structure. However, as is shown in Figure 3, as the kernel size changes, the performance of Mamba4Rec becomes quite unstable compared with our GLINT-RU. Furthermore, our GLINT-RU consistently outperforms Mamba4Rec across all kernel sizes, further demonstrating the stability and superiority of this novel GLINT-RU framework.
The hist plot in Figure 3 shows the GPU memory occupation with different kernel sizes. Setting different kernel sizes for the model does not significantly increase GPU utilization, indicating that the appropriate kernel size can be freely chosen based on accuracy requirements. Moreover, increasing the kernel size has slight impacts on the training and inference time of each mini-batch, which further verifies the efficiency and stability of our model.

### v2

@@ 3.6. Complexity Analysis
In this subsection, we will explain why GLINT-RU has inherent superiority over other popular SRS models in model efficiency.
Given that the sequence length is , the embedding size is and the kernel size for GLINT-RU is , the time complexity of GLINT-RU is . The complexity is calculated throughout the network, from the embedding layer to the prediction layer. Discussion. Our GLINT-RU shows significantly low and linear time complexity, as GLINT-RU is a highly paralleled mixed network with only one layer to achieve high performance. Our framework utilizes paralleled expert networks and employs an efficient GRU module to capture long-term dependencies. GLINT-RU is more efficient than other models in the following aspects:
GLINT-RU v.s. Transformer-based SRSs: Firstly, traditional transformer-based model (Kang and McAuley, 2018) suffers from large computational complexity, especially when the sequence length is large. Linear attention mechanism (Liu et al., 2023a) can be applied to substantially reduce the computational cost. However, they still require stacked transformer structures to achieve high performance, while GLINT-RU achieves outstanding performance with only one layer.
GLINT-RU v.s. MLP-based SRSs: Secondly, The inference speed of the MLP-based models (Gao et al., 2024) may decrease when faced with long sequence length, as the sequence mixing layer has quadratic time complexity. In contrast, the paralleled expert module of GLINT-RU is more suitable for processing long sequential data.
GLINT-RU v.s. SSM-based SRSs: Thirdly, state-space models (Gu and Dao, 2023) may require complex matrix operations and recursive calculations, which may be difficult to parallelize efficiently in practical calculations. In contrast, GLINT-RU’s paralleled expert module has a simpler update mechanism and higher efficiency.
@@ 4. Experiments
In this section, we conduct extensive experiments to show the effectiveness and efficiency of our GLINT-RU Framework. After we introduce our implementation details, the experiment results will be analyzed in detail. The experiments in this section are set to answer the following research questions:
@@
RQ3: How do the dense selective GRU, the linear attention mechanism, and the gated MLP contribute to GLINT-RU?
@@ 4.1. Datasets and Evaluation Metrics
We evaluate GLINT-RU based on three benchmark datasets ML-1M 222https://grouplens.org/datasets/movielens/, Amazon-Beauty and Amazon video Games 333https://cseweb.ucsd.edu/ jmcauley/datasets.html#amazon_reviews. Below is the basic introduction of MovieLens-1M, Amazon Beau-ty, and Amazon Video Games datasets.
MovieLens 1M: comprises user ratings of movies. It includes about 1 million anonymous ratings from users who joined MovieLens. The dataset provides information about user IDs, movie IDs, ratings, and timestamps. It is widely used for research in collaborative filtering and recommendation systems.
Amazon Beauty: is a subset of the Amazon product data, focusing specifically on beauty products. It includes user reviews and ratings of various beauty products available on Amazon. The dataset contains metadata such as product descriptions, categories, prices, and brand information. It is valuable for research in sentiment analysis, product recommendations, and consumer behavior analysis within the beauty industry.
Amazon Video Games: is a subset of the Amazon product data, focusing specifically on video game products. It includes user reviews and ratings of various video game products available on Amazon. The dataset contains metadata such as product descriptions, categories, price, and platform information. It is also valuable for research in sentiment analysis, product recommendation, and consumer behavior analysis in the video games domain.
Statistical information of the three datasets is shown in Table 1, where all the three datasets are sparse. The two Amazon datasets have relatively small data volumes, about 200,000, but they have a large number of users and items, and the average length of interaction sequences is short. In contrast, ML-1M has a larger data volume, reaching 1 million, but the number of users and items is small, with much longer interaction sequence lengths.
We adopt Recall, Mean Reciprocal Rank (MRR), and Normalized Discounted Cumulative Gain (NDCG) as the evaluation metrics for our experiments. The interactions are grouped by users chronologically, and the datasets are split by the leave-one-out strategy (Liu et al., 2023a). To be more specific, the penultimate item of the interaction sequence is used for validation. Therefore, the size of validation set is determined by the number of users.
@@ 4.2. Baselines
In this paper, we compare our GLINT-RU with two types of baseline models, i.e., traditional SRS models and efficient SRS models. The traditional SRS model includes various SRS benchmarks, while the efficient SRS models improve existing model’s structure and computational methods, thus significantly enhancing the model efficiency. Adopted baselines are listed as follows:
Traditonal SRS models
(1) GRU4Rec (Hidasi et al., 2015): utilizes GRUs to capture sequential dependencies within user interactions for session-based recommendations.
(2) BERT4Rec (Sun et al., 2019): adapts the Bidirectional Encoder Representations from Transformers (BERT) architecture to model user behaviors for personalized recommendation.
(3) SASRec (Kang and McAuley, 2018): captures long-term and short-term user preferences by applying a multi-head attention mechanism.
Efficient SRS Models
(1) LinRec (Liu et al., 2023a): reduces the computational costs substantially by changing the dot product of attention mechanism in the transformer-based models. We select SASRec as the backbone of LinRec.
(2) SMLP4Rec (Gao et al., 2024) uses a tri-directional fusion scheme to learn correlations on sequence, channel, and feature dimensions efficiently.
(3) LightSAN (Fan et al., 2021) projects the initial interactions into representations with shorter lengths, which is also an efficient approach for transformer-based models.
(4) Mamba4Rec (Liu et al., 2024a): explores the potential of selective SSMs for efficient sequential recommendation, which substantially improve the efficiency of SRS models.
@@
| Models | ML-1M | Amazon Beauty | Amazon Video Games | ||||||
@@ 4.3. Implementation
In this subsection, we introduce the implementation details of the GLINT-RU. We use Adam optimizer (Kingma and Ba, 2014) with the learning rate of 0.001 for our training process. Both the train and evaluation batch size are set as 2048. The hidden size is set as 128 for ML-1M and 64 for Amazon Beauty and Video Games. As is shown in Table 1, the average length of ML-1M, Beauty, and Video Games are 165, 8.88, and 9.54, so we set the maximum sequence length as 200 for ML-1M, and 100 for the two amazon datasets. We adopt the dropout rate of 0.5 for Amazon datasets considering their high level of sparsity, compared with 0.2 for ML-1M. We construct the attention-based baselines with 2 transformer layers so that they can achieve high performance. Other implementation details follow the settings of original papers (Liu et al., 2024a; Gao et al., 2024; Liu et al., 2023a). Please find the complete hyperparameter settings and more implementation details in Appendix B.
@@ 4.4. Overall Performance (RQ1)
In this subsection, we compare the performance of GLINT-RU with both traditional recommendation frameworks and state-of-the-art efficient models. The results, as shown in Table 2, demonstrate the effectiveness of GLINT-RU on metrics Recall@10, MRR@10, and NDCG@10 in bold. According to the above table, it is evident that GLINT-RU defeats all the selected transformer-, RNN-, MLP-, and SSM-based baselines. We improve the performance upper bound of efficient recommendation models by .
Traditional RNN-based models, like GRU4Rec, might have difficulty in dealing with complex user behaviors. Although it can capture the long-term dependencies from long sequences, it struggles to learn effectively from extremely sparse datasets. Traditional attention-based methods like BERT4Rec and SASRec have great performance on the three datasets, but it is quite inefficient due to the high computational complexity of the attention mechanism.
Efficient model LinRec changes the softmax operation, and takes attention scores from more positions into consideration, improving the performance on long-term sequential recommendation tasks compared with its backbone SASRec. SMLP4Rec and Mamba4Rec achieve impressive performance on the three datasets. However, Mamba4Rec shows enhanced proficiency in modeling long sequences (ML-1M) but exhibits low performance in relatively short sequences (Beauty and Video Games). Conversely, the SMLP4Rec shows superior performance in tasks with short sequences while being less effective with longer sequences. Additionally, SMLP4Rec requires features from the items to enhance its performance. Uniquely, GLINT-RU integrates the advantages of a linear attention mechanism and dense selective GRU module, adaptively extracting dependencies from recurrent latent item representations, fine-grained positional representations, and important semantic features. The gate mechanism employed in GLINT-RU substantially enhances its ability to filter the information based on the dynamic data environment and mix the experts according to the data adaptively.
@@
| ML-1M | BERT4Rec | 88ms | 285s/epoch | 21.51GB |
| SASRec | 55ms | 172s/epoch | 21.51GB | |
| LinRec | 37ms | 101s/epoch | 11.67GB | |
| LightSANs | 43ms | 130s/epoch | 16.99GB | |
| SMLP4Rec | 51ms | 151s/epoch | 16.13GB | |
@@
| GLINT-RU | 31ms | 86s/epoch | 8.81G | |
@@
| Beauty | BERT4Rec | 1372ms | 13s/epoch | 11.69GB |
| SASRec | 444ms | 8.1s/epoch | 7.67GB | |
| LinRec | 340ms | 5.6s/epoch | 4.14GB | |
| LightSANs | 427ms | 7.2s/epoch | 4.57GB | |
| SMLP4Rec | 361ms | 5.3s/epoch | 2.95GB | |
@@
| GLINT-RU | 278ms | 3.8s/epoch | 2.62G | |
@@
| Video Games | BERT4Rec | 1290ms | 15s/epoch | 10.98GB |
| SASRec | 406ms | 9.6s/epoch | 7.65GB | |
| LinRec | 327ms | 6.6s/epoch | 4.13GB | |
| LightSANs | 369ms | 8.3s/epoch | 4.55GB | |
| SMLP4Rec | 389ms | 9.1s/epoch | 3.38GB | |
@@
| GLINT-RU | 247ms | 4.5s/epoch | 2.49G | |
@@
Inference time of each mini-batch (batch size = 2048), training time and GPU memory of GLINT-RU and other baseline models. The best results are bolded, and the second best results are underlined.
@@ 4.5. Efficiency Comparison (RQ2)
In this subsection, we analyze the efficiency of GLINT-RU and state-of-the-art sequential recommendation models. We evaluate the model efficiency according to the inference time of each mini-batch, training time, and GPU memory occupation.
The results, shown in Table 3, provide several valuable insights: Firstly, by utilizing the efficient Dense GRU module and linear attention module, GLINT-RU dramatically reduces the training time and inference time, improving the training inference time by compared with most efficient recommendation baseline models. In addition, due to the low computational cost of GRU and linear attention mechanism, GLINT-RU exhibits minimal GPU memory consumption, which is comparable to the state-of-the-art SSM-based efficient model Mamba4Rec. In Section 3.6, we demonstrate that the GLINT-RU exhibits low theoretical computational complexity, as we employ the parallel networks and efficient GRU as core components of our recommender system. The theoretical analysis has been verified by the results in Table 3.
@@
4.6. Ablation Study (RQ3)
In the ablation study, we remove the gated MLP block, linear attention expert, dense selective GRU expert, and the two temporal convolution layers individually to verify the efficacy of each component. We conduct the ablation study on the ML-1M dataset, and the results are outlined in Table 4, providing insightful observations.
The results verify the essential role of the dense selective GRU module, as the performance of the model will dramatically decrease without the GRU module. This reveals the insights that the gated GRU module effectively captures the dependencies of interactions with fine-grained positional representations. The linear attention mechanism understands the interactions of relevant items in the sequence. As is shown in Table 4, it improves the performance of GLINT-RU to some extent. Adding a temporal convolution layer incorporates context information from adjacent items, resulting in an enhancement in model performance. In addition, the gated MLP block plays a similar role as the feed-forward network in our framework, which filters complex information from the expert mixing block. It is noteworthy that even without the gated MLP block our framework still outperforms all the state-of-the art efficient models, demonstrating its inherent remarkable superiority for sequential recommendation tasks. After we remove the gated MLP, the GPU memory occupation of GLINT-RU becomes 7.63GB, less than Mamba4Rec shown in Table 3, and the inference time will be reduced to 241ms. We name this framework “Light GLINT-RU”, which is more applicable to resource-constrained scenarios. We further discuss the ablation study on activation functions in Appendix C, where we highlight their impact on performance. Additionally, we provide a detailed analysis of the ablation study conducted on the Amazon Beauty and Amazon Video Games datasets in Appendix E, emphasizing the contributions of each component to the overall performance of GLINT-RU.
@@ 4.7. Parameter Analysis
We conduct parameter analysis on the dataset Amazon Beauty. We will first analyze the impact of the crucial hyperparameter kernel size in GLINT-RU, and then we will analyze the model performance as the hidden size and number of GLINT-RU layers changes. The discussion on these parameters provides valuable insights into the superiority of GLINT-RU.
Figure 3.(a)-(c) displays the model performance of GLINT-RU with different kernel sizes. Our model exhibits stable and high performance, providing a wide range of kernel sizes. This finding indicates the robustness of the GLINT-RU framework. This enhancement in performance can be attributed to the fact that a larger kernel size aggregates information from more items, thereby learning a more extensive context into the hidden state. However, as the kernel size continues to expand, dense selective GRU might incorporate irrelevant data into the output state, which might lead to a marginal decline in the accuracy. Mamba4Rec as a novel efficient SSM-based model, also employs a temporal convolution layer in the model structure. As the kernel size changes, the performance of Mamba4Rec becomes quite unstable compared with our GLINT-RU. Our GLINT-RU consistently outperforms Mamba4Rec across all kernel sizes, further demonstrating the superiority of this novel GLINT-RU framework.
As can be seen in the hist plots in Figure 3.(d) and Figure 4, increasing the kernel size has slight impacts on the training/inference time of each mini-batch and the GPU memory occupation, which further verifies the efficiency and stability of our model.
Hidden size . We change the model size by using different hidden sizes and compare the performance of GLINT-RU with state-of-the-art baselines. The results shown in Figure 5.(a)-(c) indicate that GLINT-RU achieves state-of-the-art performance at small hidden sizes. When is set as 64, the predictive efficacy of the GLINT-RU substantially surpasses the upper bounds of performance achieved by other baseline models. Achieving excellent results with a relatively small model dimension illustrates the superior expressiveness and significant advantages of the GLINT-RU. SMLP4Rec requires additional features to enhance the model accuracy, and its inference speed is significantly affected by the variation in the hidden size, demonstrating relatively low efficiency. The prediction accuracy of Mamba4Rec is inferior and decreases dramatically as the hidden size gets larger, indicating that it struggles to learn effective information from short behavior sequences. Attention-based models SASRec, LightSAN, and LinRec show more stable results, but they can only capture item interactions to predict the ratings and require stacked transformers to achieve relatively high performance, which has negative impacts on model efficiency in Figure 5.(d).
Number of Layers . We increase the number of GLINT-RU layers from 1 to 4 and observe the model performance and efficiency. The results and the experimental details are illustrated in Appendix D. The results indicate that our single-layer GLINT-RU can achieve high model efficiency and accuracy simultaneously.
In summary, GLINT-RU effectively combines the long/short-term dependencies with effective positional representations and important interactions which enables it to be an accurate and efficient SRS under a broad range of parameter choices.